# Nonparametric Estimation of Matching Efficiency and Elasticity on a Private On-the-Job Search Platform: Evidence from Japan, 2014-2024

### v1

@@ Abstract
I use proprietary data from the online job scouting platform BizReach in Japan, spanning from 2014 to 2024, to estimate the matching function for high-skill employed workers on a private on-the-job search platform, employing a novel nonparametric approach developed by Lange and Papageorgiou (2020). This analysis is compared to the public off-the-job search platform, Hello Work. The results indicate that matching efficiency on the private platform is both more volatile and higher than that of the public platform, suggesting the increasing popularity of the private platform. The matching elasticity with respect to users consistently hovers around 0.75, while the elasticity with respect to vacancies reaches approximately 1.0, indicating a higher and more balanced elasticity compared to the Hello Work platform. Additionally, the study reveals evidence of industry-level heterogeneity on the private platform.
@@ 1 Introduction
On-the-job search plays a crucial role in labor reallocation, leading to wage and productivity improvements (Moscarini and Postel-Vinay 2017). In the U.S., job-to-job transitions account for one-third to one-half of all hires (Faberman et al. 2022). In contrast, Japan’s labor market has historically been characterized by long-term employment stability, with workers typically staying with a single employer for most of their careers. The proportion of employed individuals who changed jobs remains low (3.25 million), representing 4.8% of all employed workers in 2023. However, reflecting the Japanese government’s policies aimed at promoting job mobility, the number of employed individuals seeking to change jobs or explore new opportunities reached 10.35 million in 2023, representing 15.3% of the employed workforce. This marks the tenth consecutive period of growth and is the highest figure on record. [^1] According to the Labor Force Statistics Office of the Statistics Bureau, Ministry of Internal Affairs and Communications. Despite the growing importance of on-the-job search in understanding labor market dynamics in many countries—particularly through private platforms—empirical evidence on the extent, matching efficiency, and elasticity of this process remains limited. This contrasts sharply with the more abundant research on job search behavior among unemployed individuals.
This paper seeks to address this gap by providing new evidence on on-the-job search platforms for employed, high-skill workers. Using proprietary aggregate data from BizReach, a prominent private online job scouting platform in Japan, I estimate the matching function within the context of on-the-job search and recover matching efficiency and elasticity by applying a novel nonparametric approach developed by Lange and Papageorgiou (2020). The platform allows registered job seekers to upload resumes, become active, and receive scouting messages from companies and headhunters actively searching for specialized talent. Workers on this platform can apply to posted jobs or wait to be scouted, contrasting with conventional job search platforms where workers actively apply for vacancies. As of July 2024, more than 2.58 million employed and self-employed workers in Japan seeking to change jobs or explore new opportunities had registered on the platform, indicating a reasonable degree of representativeness of on-the-job job seekers. To compare this private platform with the public sector counterpart studied in Otani (2024), I incorporate data from the “Report on Employment Service” (Shokugyo Antei Gyomu Tokei), using month-level aggregate data to analyze the trends of matching unemployed workers with full-time vacancies via the public employment platform, Hello Work. This comparison offers insights into the differing features and outcomes between private and public job search platforms in Japan.
My results highlight significant differences in matching efficiency and elasticity between the Hello Work public employment platform and a private BizReach platform from 2014 to 2024. For Hello Work, matching efficiency remained relatively stable until 2021, after which it sharply declined. The elasticity with respect to unemployment was consistently lower, ranging from 0.1 to 0.4, while the elasticity with respect to vacancies gradually increased to 1.0, indicating that changes in vacancies had a greater impact on job matching than changes in unemployment. This offers different implications from the findings of Otani (2024), which used a broader time range. On the private platform, matching efficiency was highly volatile, peaking in 2016. Elasticities on the private platform were also more variable and generally higher than those observed in Hello Work. The elasticity with respect to users consistently hovered around 0.75, while the elasticity concerning vacancies steadily increased, reaching around 1.0 by 2024. This suggests a more responsive matching process to changes in the number of users and a more balanced elasticity between users and vacancies compared to Hello Work.
Industry-level analysis on the private platform reveals that the Consulting sector exhibits higher matching efficiency and responsiveness to labor market changes, especially after 2020, in contrast to the IT and Manufacturing sectors. The findings underscore industry heterogeneity, where sectors like IT and Internet are characterized by stable efficiency, similar to Manufacturing, while the Consulting sector experiences more variable matching efficiency. These sectoral insights provide a deeper understanding of labor market dynamics on private platforms, with clear contrasts to the public Hello Work system.
Overall, this paper offers quantitative insights from proprietary aggregate data into the matching function in on-the-job search labor markets, though it should be noted that the private platform analyzed may not fully represent the broader on-the-job search labor market in Japan.
@@ 1.1 Related literature
Second, this paper relates to the Japanese labor market literature, particularly in the context of the public off-the-job search platform, Hello Work. Otani (2024) estimate matching efficiency and mismatch in Japan’s Hello Work platform between 1972 and 2024, showing a declining trend in matching efficiency consistent with decreasing job and worker finding rates. The match elasticity with respect to unemployment is found to be between 0.5 and 0.9, while the elasticity concerning vacancies ranges from 0.1 to 0.4. In comparison, Otani and Kanayama (2024) apply a similar nonparametric method to estimate matching efficiency and elasticity in a privately-operated spot-worker platform, contrasting these findings with Hello Work’s part-time data. My paper complements and extends these findings by focusing on a high-skill, employed worker platform, filling an important gap in the literature. Table 1 provides an overview of recent empirical studies on Japanese labor markets between January 2014 and April 2024, contributing new evidence on matching efficiency, elasticity, and mismatch.
Table 1: Nonparametric estimation of the matching function in Japanese Labor markets
Paper Platform operator Worker status Vacancy type
Otani (2024) Administrative Unemployed Full-time, part-time
Otani and Kanayama (2024) Timee Unemployed Part-time, spot work
This paper BizReach Employed Full-time
@@
Table 2: Literature on private online job search platforms (U.S)
Paper Period Platform Worker Vacancy
U.S.
Kuhn and Skuterud (2004) 98-00 Any U Both
Kuhn and Mansour (2014) 05-08 Any U Both
Faberman et al. (2022) 13-17 Any Both Both
Hershbein and Kahn (2018) 07,10–15 BGT U Both
Hershbein and Macaluso (2018) 07,10-17 BGT Both Both
Burke et al. (2020) 07,10-17 BGT Both Both
Adams-Prassl et al. (2024) 13-19 BGT Both Both
Schubert et al. (2024) 11-19 BGT Both Both
Bana (2021) 19-20 BGT Both Both
Greenwich.HR
Marinescu (2017) 07-11 CareerBuilder Both Both
Marinescu and Rathelot (2018) 12 CareerBuilder U Both
Marinescu and Wolthoff (2020) 12 CareerBuilder U Both
Azar et al. (2020) 16 CareerBuilder Both Both
Azar et al. (2022) 12 CareerBuilder U Both
Kroft and Pope (2014) 06 Craigslist Both Both
Sockin et al. (2021) 13-20 Glassdoor Both Both
Gadgil and Sockin (2020) 08-19 Glassdoor Both Full-time
Sockin (2022) 08-21 Glassdoor Both Both
Roussille and Scuderi (2023) NA Hired.com Both Both
Roussille (2024) NA Hired.com Both Full-time
Sinclair and Gimbel (2020) 14-19 Indeed Both Full-time
Maurya and Telang (2018) NA LinkedIn Both Both
Brenčič and McGee (2023) 06 Monster.com Both Both
Faberman and Kudlyak (2019) 10-11 SnagAJob Both, low-skill Part-time
Brown and Matsa (2016) 08-09 Private Both Both
Li et al. (2024) 16-19 Private Both, high-skill Full-time
Note: BGT=Burning Glass Technologies, U=Unemployed, E=Employed, Both=U&E
Table 3: Literature on private online job search platforms (Non US)
Paper Period Platform Worker Vacancy
Kuhn and Shen (2013) 08-10, China Zhaopin.com Both Both
Martins (2018) 12-16, Mozambique Emprego.co.mz Both Both
Banfi and Villena-Roldan (2019) 08-14, Chile trabajando.com Both Both
Banfi et al. (2019) 08-16, Chile trabajando.com Both Full-time
Adrjan and Lydon (2019) 14-19, Ireland Indeed Both Both
Kuhn et al. (2020) 10, China XMRC.com Both Full-time
Chen and Luo (2022) 18-19, China 51job.com Both Both
Arceo-Gomez et al. (2022) 18-20, Mexico OMBC Both Both
Choi et al. (2022) 10-20, Chile trabajando.com Both Both
Banfi et al. (2022) 10-19, Chile trabajando.com Both Full-time
Afridi et al. (2022) 19-21, India HelpersNearMe Both, blue Both
Kang and Shen (2022) 18-19, China XMRC.com Both Full-time
Barwick et al. (2023) 16-18, China Zhaopin.com Both Both
58.com
He et al. (2023) 23, China Private Both Both
Kiss et al. (2023) 22, South Africa SAYouth.mobi Both, low-skill Both
Brinatti et al. (2023) 19 & 20, global Private Both Both
Rouwendal and Koster (2023) 18, Netherlands Any Both Both
Grasso and Tatsiramos (2023) 17-19, Italy WollyBi Both Both
Escudero et al. (2024) 10-20, Uruguay BuscoJobs Both Both
Subramanian et al. (2024) 17-22, Pakistan Job Talash Both Both
Jensen (2024) 07-17, Denmark HBS Economics E Both
Benner1 et al. (2024) 17-22, Germany Private Both Both
Bassier et al. (2024) 17-19, UK Adzuna Both Both
Zvedelikova (2024) 18-19, Japan Doda Both Full-time
Otani and Kanayama (2024) 18-23, Japan Timee U, low-skill Part-time
This paper 14-23, Japan BizReach E, high-skill Full-time
Note: U=Unemployed, E=Employed, Both=U&E, OMBC=OOC Mundial, Bumeran, & CompuTrabajo
Third, this paper contributes to the expanding literature on online job search platforms. The analysis of job matching within real-world market institutions has gained prominence due to the increasing availability of data from online job platforms (Autor 2019), as summarized in Tables 2 and 3.333Examples include studies like Kuhn and Skuterud (2004), Kuhn and Mansour (2014), and Kroft and Pope (2014), which focus on worker status, while others, such as Kuhn and Shen (2013), Hershbein and Kahn (2018), Brown and Matsa (2016), and Azar et al. (2020), focus solely on vacancy data. Moreover, research like Banfi and Villena-Roldan (2019), Marinescu and Rathelot (2018), Marinescu and Wolthoff (2020), and Azar et al. (2022) incorporates both worker and vacancy information. Much of the literature emphasizes application-level or vacancy-level behavior to assess search behavior and wage elasticity. For instance, Faberman and Kudlyak (2019) leverage proprietary application-level data from an online job search engine to explore the relationship between search intensity and duration, primarily focusing on lower-skill, hourly jobs for employed and unemployed workers. Similarly, Kambayashi et al. (2023) estimate elasticities of application, interview attendance, and offer acceptance relative to posted wages using detailed process-level data from private job search and matching intermediary platforms in Japan, which became the most significant recruitment channel in 2023, as shown in Figure 1. In contrast, this paper adopts a broader macro-level perspective, evaluating the overall efficiency of the private matching platform. To the best of my knowledge, this is the first paper to estimate matching efficiency and elasticity in an online job scouting platform, offering relatively long-term insights into private online job search trends, complementing the micro-level studies.
@@ 3 Model
To estimate the matching function and recover matching efficiency, I adopt the novel approach proposed by Lange and Papageorgiou (2020). [^5] Lange and Papageorgiou (2020) additionally incorporate search effort (Mukoyama et al. 2018) and a recruitment intensity index (Davis et al. 2013). The paper highlights two critical issues: the endogeneity of matching efficiency (Borowczyk-Martins et al. 2013) and the overly restrictive nature of the Cobb-Douglas specification, which assumes fixed matching elasticity. To address these limitations, Lange and Papageorgiou (2020) propose a nonparametric identification and estimation framework for matching efficiency under specific conditions that will be discussed later.
@@
I identify the matching function and the unobserved, time-varying matching efficiency, $A$. First, I assume that $V$ and $A$ are conditionally independent given $U$, i.e., $A\perp V\mid U$. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ exhibits constant returns to scale (CRS). [^6] To align with the original model of Matzkin (2003), the function $H=m(AU,V)$ can be reformulated as $H/U=m(A,V/U)$ under CRS, where $H/U$ and $V/U$ are the job-finding rate and market tightness, respectively. These two assumptions are commonly used in the literature. Applying the nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) demonstrates that the joint distribution $G(H,U,V)$ identifies $F(A,U)$ and the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point, denoted as $A_{0}$, within the support of $(A,U,V)$. [^7] In Otani (2024), I report finite sample performance and extend the methodology through Monte Carlo simulation. Simulation results with a sample size of $T=50$ indicate that the sample size in this paper is sufficient for accurately recovering matching efficiency. For practical issues, see Brancaccio et al. (2020b), which applies this approach to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023).
@@ 5.1 Matching efficiency and elasticity in the platform
Figure 3 presents a comparison of matching efficiency and matching elasticity between the Hello Work public employment platform (left panel) and a private scouting platform (right panel) from 2014 to 2024. In the Hello Work platform, as depicted in panel (a), matching efficiency remains relatively stable around the baseline (normalized to January 2014) until 2021 but sharply declines thereafter. Panel (b) illustrates the matching elasticities with respect to unemployment and vacancies. The elasticity with respect to unemployment remains consistently lower, around 0.4, while the elasticity with respect to vacancies shows a gradual increase, reaching values near 1.0 by 2022. This suggests that changes in vacancies have a larger impact on the number of matches than changes in unemployment. [^9] The estimated elasticities differ from those reported in Otani (2024), largely due to the different time horizons considered. Specifically, Otani (2024) includes data from 1972 to 2024, a period that encompasses various economic booms and busts, which likely captures a broader range of labor market dynamics and affects the elasticity estimates.
On the private scouting platform (right panel), panel (a) shows significantly higher volatility in matching efficiency, especially after 2016, where it peaks sharply above 900 before declining and stabilizing around 400 by 2024. This indicates that the private platform experienced large fluctuations in its ability to efficiently match job seekers with vacancies, potentially due to shifts in demand or changes in the platform’s user base. Panel (b) displays the matching elasticities with respect to users and vacancies on the private platform. The elasticity with respect to users remains consistently around 0.75, while the elasticity with respect to vacancies is generally higher, rising steadily and reaching around 1.0 by 2024. This suggests a more responsive matching process to changes in the number of vacancies. The higher volatility and responsiveness of the private platform, along with a more balanced elasticity between users and vacancies, highlight a key distinction from Hello Work, where the dynamics are comparatively more stable but less responsive.

### v4

@@ Abstract
I analyze proprietary data from BizReach (2014–2024) to estimate the matching function for high-skill workers on a private on-the-job search platform using Lange and Papageorgiou (2020) nonparametric approach. Comparing it to Hello Work, I find that matching efficiency on the private platform is both more volatile and higher, reflecting its growing popularity. Matching elasticity with respect to users is around 0.75, while for vacancies, it reaches 1.0, suggesting a more balanced elasticity than Hello Work. The study also uncovers industry-level heterogeneity, highlighting differences in matching dynamics across sectors.
@@ 1 Introduction
On-the-job search plays a crucial role in labor reallocation, driving wage and productivity improvements (Moscarini and Postel-Vinay 2017). In the U.S., job-to-job transitions account for one-third to one-half of all hires (Faberman et al. 2022). While Japan has historically been characterized by long-term employment stability, recent trends indicate a growing prevalence of on-the-job search. The proportion of employed individuals changing jobs remains relatively low at 4.8% in 2023, but the share of workers actively seeking new opportunities reached a record high of 15.3%, marking the tenth consecutive year of growth. [^1] According to the Labor Force Statistics Office of the Statistics Bureau, Ministry of Internal Affairs and Communications. Despite the increasing significance of on-the-job search in labor markets worldwide, empirical research on its matching efficiency and elasticity—especially in the context of private job platforms—remains scarce. This stands in contrast to the more extensive literature on job search behavior among unemployed individuals.
This paper fills this gap by examining long-term trends in on-the-job search platforms for high-skill workers across multiple labor markets. Using proprietary aggregate data from BizReach, a leading private online job scouting platform, I estimate the matching function and recover matching efficiency and elasticity through a nonparametric approach developed by Lange and Papageorgiou (2020). Unlike conventional job boards where workers actively apply for vacancies, BizReach allows registered job seekers to receive scouting messages from companies and headhunters searching for specialized talent. As of July 2024, over 2.58 million employed and self-employed workers were registered on the platform, underscoring its relevance in analyzing the dynamics of on-the-job search.
To contextualize these findings, I compare the private platform’s performance with the public employment service, Hello Work, using month-level aggregate data from the “Report on Employment Service” (Shokugyo Antei Gyomu Tokei). This comparison provides broader insights into how private and public job search platforms function across different labor markets.
My results reveal key distinctions in matching efficiency and elasticity across platforms. For Hello Work, matching efficiency remained stable until 2021 before declining sharply. Elasticity with respect to unemployment was consistently low (0.1 to 0.4), whereas elasticity with respect to vacancies increased to 1.0, suggesting that vacancy fluctuations played a larger role in job matches than unemployment trends. In contrast, matching efficiency on BizReach was highly volatile, peaking in 2016, and exhibited generally higher elasticities than Hello Work. Elasticity with respect to users hovered around 0.75, while elasticity concerning vacancies rose steadily to 1.0 by 2024, indicating a more responsive matching process.
An industry-level analysis of the private platform highlights significant sectoral differences. The Consulting sector demonstrated higher matching efficiency and responsiveness to labor market fluctuations, particularly after 2020, whereas the IT and Manufacturing sectors showed more stable efficiency trends. These findings emphasize industry heterogeneity in labor market dynamics and suggest that private platforms function differently across sectors compared to public employment services.
By leveraging proprietary aggregate data, this study provides novel empirical insights into the matching function of on-the-job search labor markets beyond a single national context. While the BizReach platform serves as a valuable case study, the findings contribute to a broader understanding of private job market dynamics and their role in labor reallocation across different economies.
@@ 1.1 Related literature
Second, this paper relates to the Japanese labor market literature, particularly in the context of the public off-the-job search platform, Hello Work. [^3] See the survey of Miyamoto (2025) for an overview of labor markets in Japan. Otani (2024) estimate matching efficiency and mismatch in Japan’s Hello Work platform between 1972 and 2024, showing a declining trend in matching efficiency consistent with decreasing job and worker finding rates. The match elasticity with respect to unemployment is found to be between 0.5 and 0.9, while the elasticity concerning vacancies ranges from 0.1 to 0.4. In comparison, Kanayama and Otani (2024) apply a similar nonparametric method to estimate matching efficiency and elasticity in a privately-operated spot-worker platform, contrasting these findings with Hello Work’s part-time data. My paper complements and extends these findings by focusing on a high-skill, employed worker platform, filling an important gap in the literature.
@@
Third, this paper contributes to the expanding literature on online job search platforms. The analysis of job matching within real-world market institutions has gained prominence due to the increasing availability of data from online job platforms (Autor 2019). [^4] Examples include studies like Kuhn and Skuterud (2004), Kuhn and Mansour (2014), and Kroft and Pope (2014), which focus on worker status, while others, such as Kuhn and Shen (2013), Hershbein and Kahn (2018), Brown and Matsa (2016), and Azar et al. (2020), focus solely on vacancy data. Moreover, research like Banfi and Villena-Roldan (2019), Marinescu and Rathelot (2018), Marinescu and Wolthoff (2020), and Azar et al. (2022) incorporates both worker and vacancy information. Much of the literature emphasizes application-level or vacancy-level behavior to assess search behavior and wage elasticity. For instance, Faberman and Kudlyak (2019) leverage proprietary application-level data from an online job search engine to explore the relationship between search intensity and duration, primarily focusing on lower-skill, hourly jobs for employed and unemployed workers. Similarly, Kambayashi et al. (2025) estimate elasticities of application, interview attendance, and offer acceptance relative to posted wages using detailed process-level data from private job search and matching intermediary platforms in Japan, which became the most significant recruitment channel in 2023, as shown in Figure 1. In contrast, this paper adopts a broader macro-level perspective, evaluating the overall efficiency of the private matching platform. To the best of my knowledge, this is the first paper to estimate matching efficiency and elasticity in an online job scouting platform, offering relatively long-term insights into private online job search trends, complementing the micro-level studies.
@@ 3 Model
Note that our analysis does not aim to equate the two markets, but rather to compare system-level matching efficiency patterns within a consistent empirical framework. While the matching function itself is agnostic to the interpretation of matching efficiency, the implications differ across platforms. In the case of Hello Work, matching efficiency reflects the search effort of unemployed, typically lower-skilled workers seeking full-time jobs, or the effectiveness of Hello Work’s employment support services. In contrast, matching efficiency on the BizReach platform reflects the passive search effort of employed high-skilled workers awaiting scouting offers—driven primarily by recruiters’ hiring initiatives—or the effectiveness of the platform’s services, including its user interface and recommendation algorithms.
To estimate the matching function and recover matching efficiency, I adopt the novel approach proposed by Lange and Papageorgiou (2020). The paper highlights two critical issues: the endogeneity of matching efficiency (Borowczyk-Martins et al. 2013) and the overly restrictive nature of the Cobb-Douglas specification, which assumes fixed matching elasticity. To address these limitations, Lange and Papageorgiou (2020) propose a nonparametric identification and estimation framework for matching efficiency under specific conditions that will be discussed later.
@@
I identify the matching function and the unobserved, time-varying matching efficiency, $A$. First, I assume that $V$ and $A$ are conditionally independent given $U$, i.e., $A\perp V\mid U$ [^6] Matching models often incorporate a free entry condition, whereby firms continue to post job vacancies as long as the expected return is positive. This condition implies that the number of vacancies is endogenously determined and may depend on current matching efficiency, potentially violating the exogeneity assumption (Borowczyk-Martins et al. 2013). While Lange and Papageorgiou (2020) additionally incorporate search effort data (Mukoyama et al. 2018) and a recruitment intensity index data (Davis et al. 2013) to attempt to address this endogeneity issue using time-use survey data—which is typically unavailable in most contexts—Otani (2024) instead demonstrates through Monte Carlo simulations that the resulting bias is economically negligible, with estimated parameters deviating by less than 5% from the time-varying true values, and benefit of capturing time-varying efficiency is substantial. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ exhibits constant returns to scale (CRS). [^7] To align with the original model of Matzkin (2003), the function $H=m(AU,V)$ can be reformulated as $H/U=m(A,V/U)$ under CRS, where $H/U$ and $V/U$ are the job-finding rate and market tightness, respectively. These two assumptions are commonly used in the literature. Applying the nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) demonstrates that the joint distribution $G(H,U,V)$ identifies $F(A,U)$ and the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point, denoted as $A_{0}$, within the support of $(A,U,V)$. [^8] In Otani (2024), I report finite sample performance and extend the methodology through Monte Carlo simulation. Simulation results with a sample size of $T=50$ indicate that the sample size in this paper is sufficient for accurately recovering matching efficiency. For practical issues and application in different contexts, see Brancaccio et al. (2020b), which applies this approach to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023), and Otani (2025) which applies this approach to marriage markets.
@@ 5.1 Matching efficiency and elasticity in the platform
Figure 3 presents a comparison of matching efficiency and matching elasticity between the Hello Work public employment platform (left panel) and a private scouting platform (right panel) from 2014 to 2024. I set the reference period to $t=0$ corresponding to January 2014, and normalize matching efficiency by setting $A_{0}=100$. In the Hello Work platform, as depicted in panel (a), matching efficiency remains relatively stable around the baseline (normalized to January 2014) until 2021 but sharply declines thereafter. Panel (b) illustrates the matching elasticities with respect to unemployment and vacancies. The elasticity with respect to unemployment remains consistently lower, around 0.4, while the elasticity with respect to vacancies shows a gradual increase, reaching values near 1.0 by 2022. This suggests that changes in vacancies have a larger impact on the number of matches than changes in unemployment, and the assumption of constant returns to scale has been seemingly violated after 2020.101010The estimated elasticities differ from those reported in Otani (2024), largely due to the different time horizons considered. Specifically, Otani (2024) includes data from 1972 to 2024, a period that encompasses various economic booms and busts, which likely captures a broader range of labor market dynamics and affects the elasticity estimates.
On the private scouting platform (right panel), panel (a) shows significantly higher volatility in matching efficiency, especially after 2016, where it peaks sharply above 900 before declining and stabilizing around 400 by 2024. This indicates that the private platform experienced large fluctuations in its ability to efficiently match job seekers with vacancies, potentially due to shifts in demand or changes in the platform’s user base. Panel (b) displays the matching elasticities with respect to users and vacancies on the private platform. The elasticity with respect to users remains consistently around 0.75, while the elasticity with respect to vacancies is generally higher, rising steadily and reaching around 1.0 by 2024, and the assumption of constant returns to scale has been seemingly violated after 2020. This suggests a more responsive matching process to changes in the number of vacancies. The higher volatility and responsiveness of the private platform, along with a more balanced elasticity between users and vacancies, highlight a key distinction from Hello Work, where the dynamics are comparatively more stable but less responsive.